### v1

# PaECTER: Patent-level Representation Learning using Citation-informed Transformers

Abstract
PaECTER is a publicly available, open-source document-level encoder specific for patents. We fine-tune BERT for Patents with examiner-added citation information to generate numerical representations for patent documents. PaECTER performs better in similarity tasks than current state-of-the-art models used in the patent domain. More specifically, our model outperforms the next-best patent specific pre-trained language model (BERT for Patents) on our patent citation prediction test dataset on two different rank evaluation metrics. PaECTER predicts at least one most similar patent at a rank of 1.32 on average when compared against 25 irrelevant patents. Numerical representations generated by PaECTER from patent text can be used for downstream tasks such as classification, tracing knowledge flows, or semantic similarity search. Semantic similarity search is especially relevant in the context of prior art search for both inventors and patent examiners. PaECTER is available on Hugging Face. [^1] https://huggingface.co/mpi-inno-comp/paecter
1 Introduction
Domain-specific language models have become increasingly vital in social sciences and technology, transforming text into valuable data. Social scientists, in particular, have shown a keen interest in conducting similarity comparisons between patents. Existing state-of-the-art tools are tailored explicitly to scientific English, which may differ from patent-specific language. The PaECTER (Patent Embeddings using Citation-informed TransformERs) encodes patents into numerical representations to detect semantic similarity.
The introduction of the Bi-Directional Encoder Representations from Transformers (BERT) by Devlin et al., (2019) was a major advancement in the toolbox of scientists working with text. These kinds of models are trained on huge textual corpora and generate fixed-size dense vectors (embeddings) from text. BERT and its derivatives work with a fixed vocabulary of so-called tokens (words or sub-words). Tokens not found in this vocabulary are either discarded or fragmented into smaller, recognizable tokens. A set of tokens may not adequately encompass the complete spectrum of linguistic features found in particular language usage, leading to the creation of language-specific BERT derivatives.
Therefore, researchers have developed various BERT models that have been trained on and for specific corpora. Some of them are suitable for the task of transforming patent texts into embeddings, namely BERT for Patents (Srebrovic and Yonamine,, 2020) or SciBERT (Beltagy et al.,, 2019), a BERT-based pre-trained language model specific to scientific language. Having the right vocabulary is crucial for domain-specific tasks. Our analysis reveals that only approximately 50% of the vocabulary used in SciBERT directly overlaps with that of BERT for Patents.
However, since researchers are interested in the similarity between texts, such models have a drawback. They are not ideally suited for such a similarity task, as their training goal was to predict masked tokens and next sentences, not similar documents.
Therefore, a suitable relatedness/similarity signal and a similarity training objective must be identified. To accomplish such a task, contrastive learning is used. This machine learning method finds patterns by teaching the model to recognize similarities and differences between data points during the training. The key element is the triplet loss function and specially crafted training data that contain signals for similarity and dissimilarity.
SPECTER (Cohan et al.,, 2020)and SPECTER2 (Singh et al.,, 2022) use this approach. They make use of the existing capabilities to model scientific language by taking SciBERT as a foundation and generate a training dataset with so-called triplets. These triplets consist of a query document, an actual (positive) citation, and a non-cited (negative) document. The SPECTER model initially used 684,000 training triplets (146,000 query documents), while its extended version, SPECTER2, was trained on over 6 million triplets. By encoding titles and abstracts of scientific publications, SPECTER and SPECTER2 have shown improved performance on multiple downstream tasks in the scientific domain, such as topic classification, citation prediction, and recommendation.
SEARCHFORMER (Vowinckel and Hähnke,, 2023) uses the same approach in the patent domain. However, this model is unsuitable for comparing patent titles and abstracts because of its focus is on the patent claims. Furthermore, it is not publicly available.
This leaves a gap in the patent space for a model with the appropriate vocabulary/language and citation-induced similarity that is publicly available at the same time.
We fill this gap by following the contrastive learning technique used in the previously described models. The foundation of our model is the BERT for Patents, whose learning training objective is to find the semantic similarity between the cited patent and the focal patent with respect to the non-cited patent.
Since we rely on citations as similarity signal, we focus on EPO patents, for two reasons. At the EPO, references references in search reports are under the full control of examiners, mitigating any strategic bias typically associated with citations made by inventors or patent attorneys. Moreover, these references are classified with a unique emphasis on their contribution to the patent filed (Webb et al.,, 2005), a feature not present in US patents. In total, we use 300.000 patent families from the EPO, and generate five triples for each.
We have fine-tuned the SPECTER 2.0 model with our training patent dataset for benchmarking. We have named this model PAT SPECTER. Figure 1, shows how the different models relate to each other.
Figure 1: Overview of the different models
Both models, the PaECTER and the PAT SPECTER can be used for downstream tasks in the patent domain, such as topic classification, citation prediction, and recommendation.
2 Training Data
Our training data involves 300,000 English-language patent families, including applications filed with the European Patent Office (EPO) from 1985 to 2022. A patent family refers to those patent applications and grants which originate from the same set of priority filings, but seek protection in different jurisdictions.
We specifically focus on EPO patents for two critical reasons. First, references in EPO search reports are under the full control of examiners, mitigating any strategic bias typically associated with citations made by inventors or patent attorneys. Second, these patents are classified with a unique emphasis on their contribution to the patent filed (Webb et al.,, 2005), a feature not present in US patents.
2.1 Objective
The primary goal of our model is to grasp an understanding of similarity and dissimilarity based on textual content. By feeding the model examples of similar and dissimilar documents in relation to a focal document, the model autonomously learns optimal weights that minimize the loss function. Triplet margin loss learning is an effective approach in this context, where the adjustments of the weights are guided by maximizing the differences between similar and dissimilar examples.
Our training dataset comprises triplets of patents with a focal patent, a similar patent (positive), and a dissimilar patent (negative). Similar patents are those that the focal patent cites directly with a specific citation category. Dissimilar patents are those it does not cite, subject to further refinements.
The triplet margin loss function is obtained as:
$\max\{(\lVert V_{F}-V_{P}\rVert_{2}-\lVert V_{F}-V_{N}\rVert_{2}+m),0\}$ (1)
where .$\lVert.\rVert_{2}$ represents the $L_{2}$ norm distance, and $V_{F},V_{P},V_{N}$ are numerical representations for the focal patent ($P_{F}$), positive patent ($P_{P}$), and negative patent ($P_{N}$), respectively. The margin variable ($m$) ensures that the positive patent with respect to the negative patent is always closer to the focal patent by at least a certain value.
2.2 Focal Patents
We selected 300,000 patent applications as focal patents from the PATSTAT 2023 Spring version following specific criteria:
  1. 1.
    They are either EP-direct or Euro-PCT applications.
  2. 2.
    They were filed between 1985222We exclude patents filed prior to 1985 due to the ongoing transition to the EPO during that period. and 2022.
  3. 3.
    They have been assigned a Cooperative Patent Classification (CPC).
  4. 4.
    They have a minimum of two backward citations of category X/Y/I, or at least one backward citation of category X/Y/I and one of category A.
  5. 5.
    Their referenced patents collectively cite at least two other patents.
Given our focus on English language training, all patents, including focal, positive, and negative ones, must have at least an English abstract. In cases where this condition is not met, we strive to substitute it with the most fitting patent sibling from the same DOCDB family according to the following priority rule:
$\begin{split}WO>US>GB>CA>AU>DE>\\ CN>TW>KR>FR>JP\end{split}$
In total, 1,358,264 patent documents satisfy these criteria. From those, we randomly sample 300,000 documents. Finally, we split the training dataset into 85:15 ratio, such that all triplets for the same focal document remain together. The 15% fraction is reserved for validation purposes during the training process.
2.3 Positive Citations
Positive citations are selected exhaustively with replacement from the pool of eligible patent documents for each focal patent, based on specific citation categories (Table 1). Categories X, Y, I, and A are pertinent, indicating a strong similarity or relevance to the state of the art.
Table 1: Citation Categories Provided by the EPO
Category Meaning
X
Particularly relevant if taken alone - prejudicing novelty.
I
Particularly relevant if taken alone - prejudicing inventive step.
Y
Particularly relevant if combined with another document of the same category.
A
Documents defining the state of the art and not prejudicing novelty or inventive step.
O
Non-written disclosure.
P
Intermediate document.
T
Theory or principle underlying the invention.
E
Earlier patent application, but published after the filing date of the application searched (potentially conflicting patent documents).
D
Document cited in the application.
L
Document cited for other reasons.
Member of the same patent family, corresponding document
Source From Webb et al., (2005). Categories in bold font are the ones we are using.
Of these categories, X, Y, I and A are relevant for the training since the focal patents citing the patents of categories X, Y, and I have strong similarity with the cited ones. The cited patent of kind X can even prejudice the novelty of the focal patents. The cited patent of kind A also reflects the state of the art for the focal patents. Thereby, we consider the cited patents of kinds X, Y, I, and A to be similar to the focal patents.
The sampling with replacement introduces duplicates at the level of the focal patent. Duplicates may also enter if they are cited by two included patents. The positive samples of the complete training set have 901,587 unique patent application IDs out of 1,500,000 rows. However, $60\%$ of positive patents are used only once.
2.4 Negative Citations
Negative patents, representing dissimilarity with the focal patents, are selected with a nuanced approach. First, they must come from the universe of all patents ever cited by any EPO patent, which was filed during the 1985-2022 period. Second, at least the abstract is available in English through the cited patent itself or its best substitute. Third, a CPC category must exist.
We then group the negatives into easy negatives and hard negatives:
  • Easy Negatives: Uncited patents in the same CPC class and published within five years prior to the focal patent, and not indirectly cited.
  • Hard Negatives: Patents cited by the focal patent’s backward citations but not by the focal patent itself.
The selection process is illustrated in Figure 2. The number of all potential easy negatives is 6,484,216; there was no focal patent that did not meet the requirements of at least three easy negatives.
Figure 2: Positive, easy and hard negatives in patent training selection
With the selection of the negatives, there are also some duplicates. The negative samples of the complete training set have 1,251,737 unique patent application IDs out of a total of 1,500,000.
2.5 Test Dataset
For the test dataset, we construct 1000 samples consisting of 1 focal patent with 5 positive citations and 25 negative citations. The negative citations are a random sample of all easy negatives of a patent. Since the dataset is smaller, only focal patents were selected where there were enough positive and negative examples. We do not sample with replacement.
The test dataset has no overlap with the training data. None of the focal, negative, or positive patent selections are present in the training dataset.
3 Training
We build our model utilizing sentence transformers architecture (Reimers and Gurevych,, 2019) and HuggingFace framework (Wolf et al.,, 2020). At the core of which, we use BERT for Patents (Srebrovic and Yonamine,, 2020) that is connected to an input module at the preceding layer and also connected to a document level embedding module at the following layer.
The input module concatenates a patent’s title and abstract and passes the concatenated text to the BERT for Patents model. Because of the architecture of BERT models, only the first 512 tokens (about 393 words) are used. All the 1024 dimensional tokens of the output layer of the BERT for Patents model subsequently go to the document level embedding module, which produces the mean embedding of all these tokens as the final numerical representation for the given input patent text. [^3] We experimented with the mean embedding of the tokens of the output layer of BERT for Patents model; it outperformed the embedding of [CLS] token in the similarity learning task. Hence, we do not use [CLS] token as the final representation of any given text. During training, we feed each triplet to the model. It then optimizes the triplet loss function (eq. (1)) by comparing the numerical representations of all three documents in that triplet.
We do minimal hyperparameter-tuning for training, following the leads of SPECTER and SEARCHFORMER. Post-tuning, we use learning rate of $1e$-$5$, 10% of the training data for warmup steps, and triplet loss margin ($m$) of 1 for training in addition to the optimizer, Decoupled Weight Decay Regularization (Loshchilov and Hutter,, 2019)
We train our model for 4 epochs, but the training doesn’t improve after the second epoch. Each training epoch spans approximately 20 hours, with validation occurring every 4000 steps.
We use a batch size of 4 (maximum GPU capacity) and gradient accumulation of 8 on four NVIDIA A100-SXM4-40GB GPUs for faster training and achieving an effective batch size of 128. To train the model with multiple GPUs, we have adapted the code base of the sentence transformers package slightly [^4] See the forked GitHub repository:
https://github.com/max-planck-innovation-competition/sentence-transformers
, and also utilized Hugging Face’s Accelerate library.
4 Evaluation
4.1 Performance Evaluation
We conduct a rank-aware evaluation of our trained model, PaECTER and compare it against several other models and algorithms on our test dataset. The baseline of our evaluation is the BM25 algorithm. We then compare the performance of SPECTER, SPECTER version 2 (BASE), and BERT for Patents. [^5] We do not include SEARCHFORMER because the model is not available publicly.
Furthermore, we also fine-tuned a variant of SPECTER 2 (BASE) on our patent training data and refer to it as PAT SPECTER. The primary objective of including PAT SPECTER in the comparison is to benchmark PaECTER against a model originating from scientific domain but fine-tuned on our specific domain of interest.
We use three evaluation metrics. The first Rank First Relevant (RFR), which is simply the position of the first actual reference patent within the ranked list of related patents for each sample. Secondly, we use the Mean Average Precision (MAP) assesses the average precision of correctly identifying the 5 actual reference patents within the ranked list of 30 candidate patents for each sample, and then averages these precision scores across all samples. Finally, we use Mean Reciprocal Rank @ 10 (MRR@10), which averages the reciprocal ranks of the first correctly predicted patent among the set of 10 out of 30 best ranked patents, averaged over all test samples.
Moreover, since pooling strategies can potentially affect model’s performance, we consider two widely used pooling strategy during evaluation, such as CLS and mean. Table 2 presents the performances of all models under consideration for both the pooling strategies.
In Table 2, all three metrics point to the superiority of the PaECTER model regardless of which pooling strategy between CLS and mean is used. In particular, our trained model with mean pooling achieves higher accuracy over BERT for Patents by $7.85$ points (significant, p<.001) in MAP and $7.76$ points (significant, p<.001) in MRR@10. Furthermore, Figure 3 illustrates that in terms of RFR score, our model consistently predicts first relevant one within the topmost positions of the predicted outcome compared to other models.
Table 2: Rank-aware Evaluation of Several Models
Avg. RFR MAP MRR@10
Model CLS Mean CLS Mean CLS Mean
BM25 1.79 51.84 76.61
BERT-large-uncased 2.29 1.98 38.17 51.69 58.95 71.06
SciBERT-scivocab-uncased 3.07 1.91 37.76 51.31 58.80 73.45
SPECTER 1.74 1.70 55.86 56.10 77.04 76.40
BERT for Patents 2.44 1.55 44.44 60.32 67.14 80.49
SPECTER 2.0 1.79 1.68 56.12 56.96 75.47 77.96
PAT SPECTER 1.42 1.55 63.42 60.63 84.07 80.75
PaECTER 1.32 1.31 66.91 68.17 88.03 88.25
Note: We show the performance of PaECTER model in terms of Rank First Relevant (RFR), Mean Average Precision (MAP) and Mean Reciprocal Rank (MRR) at k=10, and compare the same with other models. MAP metric uses 5 positive and 25 easy negative patents. RFR and MRR@10 use just the first positive. SPECTER and SPECTER 2.0 are pre-trained models for scientific texts. PAT SPECTER is a SPECTER 2.0 model fine-tuned on our training data. PaECTER is BERT for Patents model fine-tuned on our training data. Scores are rounded off to two decimal places. Bold values represent the optimum in each column.
Figure 3: ECDF for the distribution of RFR scores across different models
Note: The distribution of Rank First Relevant (RFR) scores for PaECTER and other models used in comparison. All models use their best pooling strategy, as determined from Table 2.
5 Conclusion
The PaECTER model can be a useful tool for researchers working with patent text. For example, researchers often rely on explicit citations to determine relatedness or knowledge flows. Not only are citations rare, but they can also be selected in a strategic or opportunistic manner (Jaffe and de Rassenfosse,, 2017).
In response to the limitations of citation-based analysis, researchers increasingly turn to textual analysis to determine relatedness. A notable instance in this evolving field is the work of (Kelly et al.,, 2021), who identified breakthrough patents by assessing their textual dissimilarity to the existing patent corpus, employing a TF-IDF transformation of patent texts. While informative, this method has computational constraints.
The PaECTER model addresses these challenges by enabling the establishment of patent relatedness in scenarios where citations are absent or may be strategically manipulated. Its application offers immediate and scalable benefits, making it a powerful tool for exploring patent landscapes and discerning innovative breakthroughs. This capability significantly contributes to the nuanced understanding of knowledge flows and patent ecosystem dynamics, thus broadening the horizons for patent research and analysis.
The models are publicly available on Huggingface:
6 Acknowledgements
Our work has been enabled by grant 2023/8401 from the European Patent Office under their Academic Research Programme (EPO-ARP) 2021. We also thank the team at the Max Planck Computing and Data Facility (MPCDF) for allowing us to use their high-performance computing cluster.